OpenStudy (anonymous):

Write the expression as a sum and/or difference of logarithms. Express powers as factors. log3(9x) "Log 3 of 9 x"

OpenStudy (anonymous):

@Ashleyisakitty @MathSucks42 @hartnn

OpenStudy (anonymous):

Please help!

hartnn (hartnn):

its \(\log_39x\) right ? lets first use the easier property, \(\log_ab\times c=\log_ab+\log_ac\) so what about \(\log_39\times x\)

OpenStudy (anonymous):

My mom is gone and I'm doing her college test for her! SHE NEEDS THIS OR SHE FAILS! PLEASE, I DONT KNOW ANY OF IT AND I HAVE TO DO THIS WHILE SHE IS AT WORK! PLEASE PLEASE :((((

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

Hartnn, I don't understand any of it. It's 11 and she has to have it done by 12. This is NOT my homework, it is hers, and if she doesn't get it, then she owes the school back a lot of money. PLEASE HELP me and my family!

OpenStudy (anonymous):

Hmmm idk, but i doubt any parent (mom or dad) would put their son or daughter to do their "college test" for them.

OpenStudy (anonymous):

If its hers, then she might have been a bit more organized and responsible to finish her school work.

OpenStudy (anonymous):

She is at work. You have to believe me. It's 11:02 and I'm freaking out because I'm only half way done. I'm a common user at OpenStudy and would NEVER pull something like this. PLEASE just help! :(

OpenStudy (anonymous):

Look, if you could just help explain it to me I MIGHT get it. But I'm really having trouble. I'm only in Algebra I

OpenStudy (anonymous):

So how do you have half of it done?

OpenStudy (anonymous):

Because I've been using Mathaway and Yahoo Answers, and some of it I can sort of understand

OpenStudy (anonymous):

Wolfram isn't helping with this one!

OpenStudy (anonymous):

Well you doing this type of math that you have not learned and do not understand, and thus asking for help, it wont help because you don't know what we are talking about, nor the terms. Yes, its great that you might want to learn, but this way is not very clever; plus its your moms test.

OpenStudy (anonymous):

Please don't do this to me :( I'm really desperate. @hartnn , could you just try to explain it again, please!!!

OpenStudy (anonymous):

I would say that you are capable of doing this, and it is actually your test.

OpenStudy (anonymous):

Just to point out, you are not much of a "common OS user" .

hartnn (hartnn):

ok, simply plug in a=9, b=x \(\log_a(b\times c)=\log_ab+\log_ac\) \(\log_3(9\times x)=...?\)

hartnn (hartnn):

***b=9,c=x

OpenStudy (anonymous):

It is my mom's. Here is a picture! http://prntscr.com/1g2jb0 Look, can we stop arguing and just try to explain.

OpenStudy (anonymous):

so log9(x * c) = log9 + x + log9^c ?? Is that it? I'm so confused.

hartnn (hartnn):